I am posting two questions, I am not getting the repeated squares method. It has been a long time that I worked this method. I am confusing myself, please help.

a. 2017^9521 mod 4724

b. 971^321 mod 765

Thank you much
Answer #1

a. 2012 9521 mod 4724 we use sulle is theotem to solve llie problems @aless theatemi: this states that if acudii ase platil2012236° = [ (ued 4224) . 20/4560X4 = 14 (mod 4424) . 201294 2. / (ived 4224). 20/24x 2012 = 7X 2014(mod 4224) : .. 201295299 Sat, mod 765 - 94 = 206 mod 765 92, 321 3 206 321 mod 265 = ((206)20 x 206 (mod: 7652 3 (361)°° x 206 (mod 765) W ood (20